OpenStudy (anonymous):

How do u graph 7x+y=5

OpenStudy (dmndlife24):

First, you have to change it to slope-int. form which is y = mx + b, m = slope, b = y-int. To do this you will have to subtract 7x from both sides... y = -7x + 5 It's simple. 5 is the y-int. and -7 is the slope.

OpenStudy (anonymous):

7x/7=5/7 1y/1=5/1 x=1.4 y=5